# A faster algorithm for Vertex Cover parameterized by solution size

### v6

@@ 8.3 Line 3: Branching on a 4-vertex with a 3-neighbor
Figure 6: We show the degrees of each vertex. In $G-u$, the 2-vertex $v_{1}$ has neighbors $x_{1},x_{2}$, which get merged into $z_{1}$. The new vertex has 4 neighbors in total and at least two 3-neighbors.
@@ 8.4 Line 5: Branching on a vertex in a 4-regular graph
We select a vertex $u$ and an ordering of its neighbors $v_{1},v_{2},v_{3},v_{4}$; our general strategy is to split on $v_{1},v_{2}$. The graph $H=G-\{v_{1},v_{2}\}-N[v_{3},v_{4}]$ plays a central role in the analysis. We let $I$ be a min-set of $H$; for brevity, we define $A=\{v_{1},v_{2}\}$ in this section.
@@
Suppose $G$ has a 6-cycle $x_{1},\dots,x_{6}$. Then $R(x_{1},N[x_{4}])\geq 2^{*}$; furthermore, either $\operatorname{codeg}(x_{1},x_{4})=1$ or $x_{4}$ is a special 4-vertex in $G-N[x_{1}]$.
@@
At this point, since $\operatorname{girth}(G)\geq 5$ and $G$ is 4-regular, each vertex $v$ has twelve vertices at distance two which all have degree 3 in $G-N[v]$. The graph $G^{\prime}=G-x_{1}-N[x_{4}]$ has two 2-vertices $x_{2},x_{6}$. We have $x_{2}\not\sim x_{6}$, as otherwise would have a triangle $x_{1},x_{2},x_{6}$. Since $G^{\prime}$ has at least 11 subcubic vertices, Proposition 34 gives $S(G^{\prime})\geq 2^{*}$.
Now suppose $\operatorname{codeg}(x_{1},x_{4})=0$. Then the graph $G-N[x_{1}]$ retains vertex $x_{4}$ and its neighbors, while $x_{3},x_{5}$ lose their neighbors $x_{2},x_{6}$ respectively. Since $\operatorname{girth}(G)\geq 5$, we have $x_{3}\neq x_{5}$ and the degrees of $x_{3},x_{5}$ are not reduced by two. So $x_{4}$ has two 3-neighbors $x_{3},x_{5}$ in $G-N[x_{1}]$. ∎
@@
  • If $\deg_{H}(x)=1$, then $x\sim A$ and $R(u,N[x])\geq 2^{*}$.
  • If $\deg_{H}(x)=2$ and $x\not\sim A$, then $x$ is a special 4-vertex in $G-N[u]$.
@@
Because $\operatorname{girth}(G)\geq 5$, we have $\operatorname{codeg}(x,v_{3})\leq 1,\operatorname{codeg}(x,v_{4})\leq 1$ and $|N(x)\cap A|\leq 1$. If $\deg_{H}(x)=1$, then necessarily $x$ has a neighbor $v_{i}\in A$ and a neighbor $r_{3}\sim v_{3}$ and a neighbor $r_{4}\sim v_{4}$ and thus $u,v_{3},r_{3},x,r_{4},v_{4}$ is a 6-cycle; by Proposition 41 we have $R(u,N[x])\geq 2^{*}$. Likewise, if $\deg_{H}(x)=2$ and $x\not\sim A$, then $x$ has neighbors $r_{3}\sim v_{3},r_{4}\sim v_{4}$, and then $G$ has a 6-cycle $u,v_{3},r_{3},x,r_{4},v_{4}$. Furthermore, $x\not\sim A$ by hypothesis and $x\not\sim\{v_{3},v_{4}\}$ since $x\in H$. Hence $\operatorname{codeg}(u,x)=0$ and so by Proposition 41 the vertex $x$ is special in $G-N[u]$. ∎
@@
If $\operatorname{minsurp}(H)<0$ and $\deg_{H}(x)\leq 2$ for all $x\in I$, then $G-N[u]$ is special.
@@
We assume that $x\sim A$ for all $x\in I$, as otherwise the result follows from Proposition 42. Since $\operatorname{mindeg}(H)\geq 1$, and $|N(I)|<|I|$ strictly, there must be some vertex $y\in H$ with two neighbors $x_{1},x_{2}\in I$. The neighbors of $x_{1},x_{2}$ in $A$ must be distinct else $G$ would have a 4-cycle. So suppose without loss of generality $x_{1}\sim v_{1},x_{2}\sim v_{2}$ and $G$ has a 6-cycle $u,v_{1},x_{1},y,x_{2},v_{2}$. Note that $y\not\sim\{v_{3},v_{4}\}$ since $y\in H$ and $y\not\sim\{v_{1},v_{2}\}$ since then $G$ would have a 3-cycle $v_{1},x_{1},y$ or $v_{2},x_{2},y$. So $\operatorname{codeg}(u,y)=0$ and hence $G-N[u]$ is special by Proposition 41. ∎
For our tie-breaking rule, we assume that if we are in a given case that ordering of neighbors $v_{1}^{\prime},\dots,v_{4}^{\prime}$ would fall into a previous case.
Case I: $\boldsymbol{\operatorname{shad}(N[v_{3}])\leq 1}$. Let $J$ be a min-set of $G-N[v_{3}]$; since $\operatorname{mindeg}(G-N[v_{3}])\geq 3$, necessarily $|J|\geq 2$, and since $G$ is in 4BP we have $\operatorname{surp}_{G-N[v_{3}]}(J)=1$. Here, we split on $v_{3}$. By Proposition 38, $G-N[v_{3}]$ has at least 8 subcubic vertices, so $R(N[v_{3}])\geq 2^{*}$ by Proposition 33. Overall, we need:
@@
Case II: $\boldsymbol{\operatorname{minsurp}(H)\leq-2}$. Since $\operatorname{mindeg}(H)\geq 1$, we must have $|I|\geq 3$. Let $J=I\cup\{v_{4}\}$. Here, $\operatorname{surp}_{G-N[v_{3}]}(J)\leq 2$, and $|J|=|I|+1\geq 4$. Since we are not in Case I, we have $\operatorname{shad}(N[v_{3}])\geq 2$ and so $J$ is a min-set in $G-N[v_{3}]$.
Here again we split on $v_{3}$ and then split on $J$ in the subproblem $G-N[v_{3}]$, generating subproblems $G-v_{3},G-N[v_{3}]-J,G-N[v_{3},J]$. Since we are not in Case II, we have $\operatorname{shad}(N[v_{3}],J)\geq\operatorname{shad}(N[v_{3}])-|J|\geq 2-|J|$ and likewise $\operatorname{shad}(N[v_{3},J])\geq\operatorname{shad}(N[v_{3}]-|N[J]|\geq 2-|N[J]|$. So the latter two subproblems have drops $(2.5,8^{*}),(2.5,10^{*})$ respectively. Here we need:
@@
Note: In all subsequent cases, our strategy is to split on $v_{1},v_{2}$, and then apply (A1) to vertex $u$ in subproblem $G-\{v_{1},v_{2}\}$. This gives four subproblems $G_{1}=G-N[u],G_{2}=G-N[v_{1}],G_{3}=G-v_{1}-N[v_{2}],H$. Here, $G_{1},G_{2},G_{3}$ have drop $(1.5,4^{*}),(1.5,4^{*}),(2,5^{*})$ directly.
@@
Note: For the remainder of the analysis, we assume $\operatorname{minsurp}(H)=-1$, and we let $r=|I|$; we have $r\geq 2$ since $\operatorname{mindeg}(H)\geq 1$. For $r\leq O(1)$, the graph $H$ has drop $(3,(8+r)^{*})$ after applying (P1).
@@
Next suppose that $r=3$. Clearly $\deg_{H}(x)\leq 2$ for all $x\in I$, so $G-N[u]$ is special by Proposition 43. Since in this case $H$ has drop $(3,11^{*})$, we then need:
@@
Finally suppose $r=2$. Again, $G-N[u]$ is special by Proposition 43. Let $I=\{x_{1},x_{2}\},N_{H}(I)=\{y\}$. Since $\deg_{H}(x_{1})=\deg_{H}(x_{2})=1$, Proposition 42 implies that $x_{1}\sim A$ and $x_{2}\sim A$; since they already have common neighbor $y$, we assume without loss of generality that $x_{1}\sim v_{1}$ and $x_{2}\sim v_{2}$. Now consider the 6-cycle $u,v_{1},x_{1},y,x_{2},v_{2}$. Here, $x_{2}$ has neighbors $v_{2},y,r_{3}\sim y_{3},r_{4}\sim y_{4}$. We must have $\operatorname{codeg}(v_{1},x_{2})=0$; for, if $v_{1}\sim y$, then $G$ would have a 3-cycle $v_{1},x_{1},y$, and if $v_{1}\sim v_{2}$, then $G$ would have a 3-cycle $u,v_{1},v_{2}$, and if $v_{1}\sim r_{i}$ for $i\in\{3,4\}$, then $G$ would have a 4-cycle $r_{i},v_{i},u,v_{1}$. So by Proposition 41, the graph $G_{2}=G-N[v_{1}]$ is also special.
Accordingly, in this case we need:
@@
Case V: $\boldsymbol{G_{3}}$ is $\boldsymbol{3^{*}}$-vul. If $\deg_{H}(x)=1$ for any $x\in I$, then by Proposition 42, $u,x$ share a neighbor $v\in A$ and $R(u,N[x])\geq 2^{*}$. But then vertex $u^{\prime}=v$ and its two neighbors $v_{1}^{\prime}=u,v_{2}^{\prime}=x$ would be covered in Case IV, or an earlier case.
Accordingly, we must have $\deg_{H}(x)\geq 2$ for all $x\in I$. At this point, consider counting the two-element subsets of $N_{H}(I)$. Each vertex $x\in I$ contributes $\binom{\deg_{H}(x)}{2}$ such subsets and since $G$ has no four-cycles these are all distinct. So $\sum_{x\in I}\binom{\deg_{H}(x)}{2}\leq\binom{|N_{H}(I)|}{2}=\binom{r-1}{2}$. Since $\deg_{H}(x)\geq 2$ for all $x\in I$, this implies that $r\geq 5$ and, furthermore, if $r=5$ then $\deg_{H}(x)=2$ for all $x\in I$; in the latter case, by Proposition 43 the graph $G-N[u]$ is special. So depending on whether $r=5$ or $r\geq 6$, we need respectively:
@@
Case VI: Everything else. Since $G$ is in 4BP, we have $\operatorname{shad}(v_{1},N[v_{3}])\geq\operatorname{shad}(N[v_{3}])-1\geq 0$. On the other hand, $\operatorname{surp}_{G-v_{1}-N[v_{3}]}(I\cup\{v_{4}\})\leq\operatorname{minsurp}(H)+1+(\deg_{G-N[v_{3}]}(v_{4})-1)=2$. So $G-v_{1}-N[v_{3}]$ is $3^{*}$-vul via indset $I\cup\{v_{4}\}$; it would be covered by Case V. So Case VI is impossible.
@@ 9 Branching on 5-vertices
There is a major difference between the degree-4 and degree-5 settings. In the former, we can ensure that the graph has many 3-vertices, and so branching on a single 4-vertex can create new 2-vertices. This structure cannot be maintained in the degree-5 case: at least two 5-vertices are jointly needed for the desired simplifications. The process is not self-sustaining.
Our strategy depends on the relative density of 5-vertices. When 5-vertices are very common, we can find closely-clustered groups of them for more advanced branching rules; the denser the 5-vertices, the more favorable the branching. We will actually develop three separate degree-5 algorithms with different “intensity” levels.
@@
Simplify $G$ If $\operatorname{maxdeg}(G)\leq 4$, then run either the MaxIS-4 algorithm or Branch4 Apply Section 5 to an arbitrary vertex of degree at least $5$.
Algorithm 6 Function Branch5-1$(G,k)$
@@
Given the potential branch-seqs from Section 5, we need the following constraints:
@@
The next algorithms Branch5-2 and Branch5-3 are targeted to medium and high density of 5-vertices respectively. They share some branching rules, so we describe them in a single listing.
1 Apply miscellaneous branching rules to ensure $G$ has nice structure.                                (We call this 5-base-position (5BP).) Branch on non-adjacent 5-vertices $u,v$ with $R(u,v)\geq 1$ and $R(u,N[v])\geq 1$. Branch on vertices $u,v$ where $\deg(u)=5,\deg_{G-u}(v)\geq 4$, and $G-\{u,v\}$ has a kite. Branch on a 4-vertex with two non-adjacent 5-neighbors Branch on a 5-vertex $u$ with neighbors $v_{1},v_{2},v_{3},v_{4},v_{5}$ where $v_{1}\sim v_{2}$ and $\deg(v_{1})=\deg(v_{3})=\deg(v_{4})=\deg(v_{5})=5$. Branch on a $5$-vertex with five $5$-neighbors. if $L=2$ then
2        Branch on a 5-vertex $u$ where $G-N[u]$ is 3-vul. Branch on non-adjacent 5-vertices $u,v$ with $R(u,v)\geq 1$. Branch on a $4$-vertex which has a triangle involving a 5-vertex. Branch on a 5-vertex with fewer than three 4-neighbors.
Run the cheaper of Branch5-$(L-1)$$(G,k)$ or the MaxIS-5 algorithm
Algorithm 7 Function Branch5-$L$$(G,k)$ for $L=2,3$
@@
Branch5-2$(G,k)$ has measure $a_{52}\mu+b_{52}k$ for $a_{52}=0.437086,b_{52}=0.080751$.
Branch5-3$(G,k)$ has measure $a_{53}\mu+b_{53}k$ for $a_{53}=0.379406,b_{53}=0.097471$.

### v7

@@ 6.2 Line 3: Branching on a 4-vertex with a 3-neighbor
Figure 9: We show the degrees of each vertex. In GuG-u, the 2-vertex $v$ has neighbors $x,y$, which get merged into $z_{1}$. The new vertex has 4 neighbors in total and at least two 3-neighbors.
@@ 6.3 Line 5: Branching on a vertex in a 4-regular graph
We select a vertex $u$ and an ordering of its neighbors $v_{1},v_{2},v_{3},v_{4}$; our general strategy is to split on $v_{1},v_{2}$ so as to create a 2-vertex in the resulting graph $G-\{v_{1},v_{2}\}$. To analyze this process, we use the following notations in Section 6.3:
$H=G-\{v_{1},v_{2}\}-N[v_{3},v_{4}],\qquad A=\{v_{1},v_{2}\}$
$I=\text{some chosen min-set of $H$},\qquad r=|I|$
The graph $H$ and its properties are critical to this analysis. We begin with a few observations.
@@
Suppose $G$ has a 6-cycle $x_{1},\dots,x_{6}$. Then $S(x_{1},N[x_{4}])\geq 2^{*}$; furthermore, either $\operatorname{codeg}(x_{1},x_{4})=1$ or $x_{4}$ is a special 4-vertex in $G-N[x_{1}]$.
@@
Since $\operatorname{girth}(G)\geq 5$, the graph $G^{\prime}=G-x_{1}-N[x_{4}]$ has two non-adjacent 2-vertices $x_{2},x_{6}$. Since $G$ is 4-regular, $G^{\prime}$ has at least 11 subcubic (the 12 vertices at distance two to $x_{4}$, minus $x_{1}$). So Proposition 33 gives $S(G^{\prime})\geq 2^{*}$. Now suppose $\operatorname{codeg}(x_{1},x_{4})=0$. Then the graph $G-N[x_{1}]$ retains vertex $x_{4}$ and its neighbors, while $x_{3},x_{5}$ lose their neighbors $x_{2},x_{6}$ respectively. Since $\operatorname{girth}(G)\geq 5$, we have $x_{3}\neq x_{5}$ and the degrees of $x_{3},x_{5}$ are not reduced by two. So $x_{4}$ has two 3-neighbors $x_{3},x_{5}$ in $G-N[x_{1}]$. ∎
@@
  • If $\deg_{H}(x)=1$, then $x\sim A$ and $S(u,N[x])\geq 2^{*}$.
  • If $\deg_{H}(x)=2$, then either $x\sim A$ or $x$ is a special 4-vertex in GN[u]G-N[u].
@@
Because $\operatorname{girth}(G)\geq 5$, we have $\operatorname{codeg}(x,v_{3})\leq 1,\operatorname{codeg}(x,v_{4})\leq 1$ and $|N(x)\cap A|\leq 1$. If $\deg_{H}(x)=1$, then necessarily $x$ has a neighbor $v_{i}\in A$ and a neighbor $r_{3}\sim v_{3}$ and a neighbor $r_{4}\sim v_{4}$ and thus $u,v_{3},r_{3},x,r_{4},v_{4}$ is a 6-cycle; by Proposition 42 we have $S(u,N[x])\geq 2^{*}$. Likewise, if $\deg_{H}(x)=2$ and $x\not\sim A$, then $x$ has neighbors $r_{3}\sim v_{3},r_{4}\sim v_{4}$, and then $G$ has a 6-cycle $u,v_{3},r_{3},x,r_{4},v_{4}$. Furthermore, $x\not\sim A$ by hypothesis and $x\not\sim\{v_{3},v_{4}\}$ since $x\in H$. Hence $\operatorname{codeg}(u,x)=0$ and so by Proposition 42 the vertex $x$ is special in GN[u]G-N[u]. ∎
@@
If $\operatorname{minsurp}(H)<0$ and $\deg_{H}(x)\leq 2$ for all $x\in I$, then GN[u]G-N[u] is special.
@@
We assume that $x\sim A$ for all $x\in I$, as otherwise the result follows from Proposition 43. Since $\operatorname{mindeg}(H)\geq 1$, and $|N(I)|<|I|$ strictly, there must be some vertex $y\in H$ with two neighbors $x_{1},x_{2}\in I$. The neighbors of $x_{1},x_{2}$ in $A$ must be distinct else $G$ would have a 4-cycle. So suppose without loss of generality $x_{1}\sim v_{1},x_{2}\sim v_{2}$ and $G$ has a 6-cycle $u,v_{1},x_{1},y,x_{2},v_{2}$. Note that $y\not\sim\{v_{3},v_{4}\}$ since $y\in H$ and $y\not\sim\{v_{1},v_{2}\}$ since then $G$ would have a 3-cycle $v_{1},x_{1},y$ or $v_{2},x_{2},y$. So $\operatorname{codeg}(u,y)=0$ and hence GN[u]G-N[u] is special by Proposition 42. ∎
We now describe the branching strategy, in particular, how to order $v_{1},\dots v_{4}$. For our tie-breaking rule, we assume that if we are in a given case that no other ordering of neighbors $v_{1}^{\prime},\dots,v_{4}^{\prime}$ would fall into a previous case.
Case I: $\boldsymbol{\operatorname{shad}(N[v_{3}])\leq 1}$. Let $J$ be a min-set of $G-N[v_{3}]$; since $\operatorname{mindeg}(G-N[v_{3}])\geq 3$, necessarily $|J|\geq 2$, and since we are after Line 1 we have $\operatorname{surp}_{G-N[v_{3}]}(J)=1$. Here, we split on $v_{3}$. By Proposition 39, $G-N[v_{3}]$ has at least 8 subcubic vertices, so $S(N[v_{3}])\geq 2^{*}$ by Proposition 32. Overall, we need:
@@
Case II: $\boldsymbol{\operatorname{minsurp}(H)\leq-2}$. Since $\operatorname{mindeg}(H)\geq 1$, we must have $r\geq 3$. Let $J=I\cup\{v_{4}\}$. Here, $\operatorname{surp}_{G-N[v_{3}]}(J)\leq 2$, and $|J|=|I|+1\geq 4$. Since we are not in Case I, we have $\operatorname{shad}(N[v_{3}])\geq 2$ and so $J$ is a min-set in $G-N[v_{3}]$.
Here we split on $v_{3}$ and in the subproblem $G-N[v_{3}]$, we apply (B-Crit) to $J$. Overall, this generates subproblems $G-v_{3},G-N[v_{3}]-J,G-N[v_{3},J]$. Since we are not in Case II, we have $\operatorname{shad}(N[v_{3}],J)\geq\operatorname{shad}(N[v_{3}])-|J|\geq 2-|J|$ and likewise $\operatorname{shad}(N[v_{3},J])\geq\operatorname{shad}(N[v_{3}])-|N[J]|\geq 2-|N[J]|$. So the latter two subproblems have drops $(2.5,8^{*}),(2.5,10^{*})$ respectively. Here we need:
@@
Note: In all subsequent cases, our strategy is to split on $v_{1},v_{2}$, and then apply (B-Pair) to vertex $u$ in subproblem $G-\{v_{1},v_{2}\}$. This gives four subproblems $G_{1}=G-N[u],G_{2}=G-N[v_{1}],G_{3}=G-v_{1}-N[v_{2}],H$ where $G_{1},G_{2},G_{3}$ have drop $(1.5,4^{*}),(1.5,4^{*}),(2,5^{*})$ directly.
@@
Note: For the remainder of the analysis, we assume $\operatorname{minsurp}(H)=-1$. We have $r\geq 2$ since $\operatorname{mindeg}(H)\geq 1$. For $r\leq O(1)$, the graph $H$ has drop $(3,(8+r)^{*})$ after applying (P1).
@@
Next suppose that $r=3$ and $H$ has drop $(3,11^{*})$. For $x\in I$, we have $\deg_{H}(x)\leq|N(I)|=r-1=2$, so GN[u]G-N[u] is special by Proposition 44. We then need:
@@
Finally suppose $r=2$ and $H$ has drop $(3,10^{*})$. Again, GN[u]G-N[u] is special by Proposition 44. Let $I=\{x_{1},x_{2}\},N_{H}(I)=\{y\}$. Since $\deg_{H}(x_{1})=\deg_{H}(x_{2})=1$, Proposition 43 implies that $x_{1}\sim A$ and $x_{2}\sim A$; since they already have common neighbor $y$, we assume without loss of generality that $x_{1}\sim v_{1}$ and $x_{2}\sim v_{2}$. Now consider the 6-cycle $u,v_{1},x_{1},y,x_{2},v_{2}$. Let us enumerate $N(x_{2})=\{v_{2},y,r_{3},r_{4}\}$ for vertices $r_{3}\sim v_{3},r_{4}\sim v_{4}$. We must have $\operatorname{codeg}(v_{1},x_{2})=0$; for, if $v_{1}\sim y$, then $G$ would have a 3-cycle $v_{1},x_{1},y$, and if $v_{1}\sim v_{2}$, then $G$ would have a 3-cycle $u,v_{1},v_{2}$, and if $v_{1}\sim r_{i}$ for $i\in\{3,4\}$, then $G$ would have a 4-cycle $r_{i},v_{i},u,v_{1}$.
So by Proposition 42, the graph $G_{2}=G-N[v_{1}]$ is also special. Accordingly, in this case we need:
@@
Case V: $\boldsymbol{G_{3}}$ is $\boldsymbol{3^{*}}$-vul. If $\deg_{H}(x)=1$ for any $x\in I$, then by Proposition 43, $u,x$ share a neighbor $v\in A$ and $S(u,N[x])\geq 2^{*}$. But then vertex $u^{\prime}=v$ and its two neighbors $v_{1}^{\prime}=u,v_{2}^{\prime}=x$ would be covered in Case IV, or an earlier case.
Accordingly, we must have $\deg_{H}(x)\geq 2$ for all $x\in I$. Now consider counting the two-element subsets of $N_{H}(I)$. Each vertex $x\in I$ contributes $\binom{\deg_{H}(x)}{2}$ such subsets and since $G$ has no four-cycles these are all distinct. So $\sum_{x\in I}\binom{\deg_{H}(x)}{2}\leq\binom{|N_{H}(I)|}{2}=\binom{r-1}{2}$. Since $\deg_{H}(x)\geq 2$ for all $x\in I$, this implies that $r\geq 5$ and, furthermore, if $r=5$ then $\deg_{H}(x)=2$ for all $x\in I$; in the latter case, by Proposition 44 the graph GN[u]G-N[u] is special. So for $r=5$ or $r\geq 6$, we need respectively:
@@
Case VI: Everything else. Since we are not in Case I, we have $\operatorname{shad}(v_{1},N[v_{3}])\geq\operatorname{shad}(N[v_{3}])-1\geq 1$. On the other hand, $\operatorname{surp}_{G-v_{1}-N[v_{3}]}(I\cup\{v_{4}\})\leq\operatorname{minsurp}(H)+1+(\deg_{G-N[v_{3}]}(v_{4})-1)=2$. So $G-v_{1}-N[v_{3}]$ is $3^{*}$-vul via indset $I\cup\{v_{4}\}$; it would be covered by Case V. Case VI is impossible.
@@ 7 Branching on 5-vertices
Our strategy for the degree-5 algorithm depends on the relative density of 5-vertices. When 5-vertices are very common, we can find closely-clustered groups of them for more advanced branching rules; the denser the 5-vertices, the more favorable the branching. We will actually develop three separate degree-5 algorithms with different “intensity” levels.
There is a major difference between the degree-4 and degree-5 algorithms. In the former, branching on a single vertex can create new 2-vertices. This cannot be maintained in the degree-5 case: at least two 5-vertices are jointly needed for the desired simplifications. The process is not self-sustaining. On the positive side, it means that we do not need to keep track of 3-vertices; as far as our analysis goes, they will be created anew at each branching step.
@@
1Simplify $G$
2If $\operatorname{maxdeg}(G)\leq 4$, then run either the MaxIS-4 algorithm or Branch4
Apply Figure 6 to an arbitrary vertex of degree at least $5$.
Algorithm 6 Function Branch5-1(G,k)(G,k)
@@
Given the potential branch-seqs from Figure 6, we need the following constraints:
@@
The next algorithms Branch5-2 and Branch5-3 are targeted to medium and high density of 5-vertices respectively. They share many branching rules, so we describe them in a single listing.
1
2Apply miscellaneous two-way branching rules to ensure $G$ has nice structure.
3Branch on non-adjacent 5-vertices $u,v$ with $S(u,v)\geq 1$ and $S(u,N[v])\geq 1$.
4Branch on vertices $u,v$ where $\deg(u)=5,\deg_{G-u}(v)\geq 4$, and $G-\{u,v\}$ has a kite.
5Branch on a 4-vertex with two non-adjacent 5-neighbors
6Branch on a 5-vertex $u$ with neighbors $v_{1},v_{2},v_{3},v_{4},v_{5}$ where $v_{1}\sim v_{2}$ and $\deg(v_{1})=\deg(v_{3})=\deg(v_{4})=\deg(v_{5})=5$.
7Branch on a 55-vertex with five 55-neighbors.
8if $L=2$ then
9   Branch on a 5-vertex $u$ where GN[u]G-N[u] is 3-vul.
10  Branch on non-adjacent 5-vertices $u,v$ with $S(u,v)\geq 1$.
11  Branch on a 44-vertex which has a triangle involving a 5-vertex.
12  Branch on a 5-vertex with fewer than three 4-neighbors.
Run the cheaper of Branch5-(L1)(L-1)$(G,k)$ or the MaxIS-5 algorithm
Algorithm 7 Function Branch5-LL(G,k)(G,k) for $L=2,3$
@@
Branch5-2(G,k)(G,k) has measure $a_{52}\mu+b_{52}k$ for $a_{52}=0.437086,b_{52}=0.080751$.
Branch5-3(G,k)(G,k) has measure $a_{53}\mu+b_{53}k$ for $a_{53}=0.379406,b_{53}=0.097471$.